OpenStudy (anonymous):

Th world record for the 100 m sprint is held by Jamaica's Usain Bolt at 9.58 (a) At this speed how long does it take bolt to run 10 m ?

OpenStudy (anonymous):

Some one please help me!!

OpenStudy (anonymous):

Assuming constant acceleration and no reaction time delay we have the equation.......> X-X0=V0(T) +1/2aT^2. V0=0 , X0=0 Therefore x=1/2aT^2.............>a=2x/T^2. a=2.18 m/s^2. So at 10 m, the time will be T=(2X/a)^1/2=3.03 seconds.

OpenStudy (anonymous):

A good answer but not quite what I wanted!